• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Huber

    Log entry time 23:47:57 on April10,2001

    Entry number 60141

    keyword=Evening Shift Summary

    16:00 - Shift starts with relatively smooth data taking after Maintenance Day completion.
    17:20-17:50 - Beam has been unstable due to Hall C fast feedback trips.
    18:10-18:45 - Beam off for "RF separation checks". As a result, beam intensity has improved from 85 to 92 uA.
    19:00 - Beam is now unstable because of injector problems. The beam trips off every few minutes, and then comes back up.
    21:10 - Beam has recently become more stable.
    23:10 - Beam off again because the inject problem has reappeared. MCC says that not only have the experts been called in, but the experts' experts are on the way.


    Run 3425 - Start=16:07 Stop=17:15 Events=1.0M
    Run 3426 - Start=17:16 Stop=19:29 Events=1.0M
    Run 3427 - Start=19:29 Stop=20:54 Events=1.0M
    Run 3428 - Start=20:55 Stop=21:54 Events=1.0M
    Run 3429 - Start=21:54 Stop=23:32 Events=946k